OpenStudy (anonymous):

can someone help me simplify this equation using identities. The equation is sin(x)/√1-sin²

OpenStudy (anonymous):

mathmale

OpenStudy (phi):

do you know an identity that uses 1 - sin^2 ?

OpenStudy (anonymous):

yeah there trigonometric identities

OpenStudy (phi):

and which one has 1 - sin^2 ?

OpenStudy (anonymous):

the very bottom one but what do I do I am very confused

OpenStudy (anonymous):

would we have to convert sin into tan or cos

OpenStudy (phi):

if it helps, a very famous identity is \[ \sin^2 x + \cos^2 x = 1\] from this you can get two useful identifies

OpenStudy (phi):

for example, if you subtract cos^2 x from both sides you get \[ \sin^2 x + \cos^2 x = 1 \\ \sin^2 x + \cos^2 x - \cos^2 x = 1 - \cos^2 x\] the left side simplifies, so you get \[ \sin^2 x = 1 - \cos^2 x\]

OpenStudy (anonymous):

then what do I do

OpenStudy (phi):

you can also subtract sin^2 x from both sides to get \[ \cos^2 x = 1 - \sin^2 x\] I would use that identity

OpenStudy (anonymous):

would we then put it into the equation and solve it so we change it to cos but how do we get rid of the square root and such

OpenStudy (phi):

the definition of \[ \sqrt{x\cdot x} = x \] no matter how complicated x is as you know \(\cos^2 x \) is short for \( \cos x \cdot \cos x \)

OpenStudy (phi):

in other words you have cos x times itself inside a square root \[ \sqrt{ \cos(x) \cdot \cos(x) } \]

OpenStudy (anonymous):

could we get rid of the square root and the squared symbol since they both will get rid of each other so what we are stuck with are 1-sin

OpenStudy (phi):

just looking at the bottom: \[ 1 - \sin^2 x \] by the identity you know that is the same as \[ \cos^2 x \] the 1- sin^2 x is *replaced* by cos^2 x

OpenStudy (phi):

so you have \[ \frac{\sin x}{\sqrt{\cos x \cdot \cos x } } \]

OpenStudy (phi):

and you know how to simplify sqr( a*a) to get a, right ?

OpenStudy (phi):

remember sqrt( anything * anything) = anything

OpenStudy (anonymous):

I get it now thanks for the help because sinx/cosx. Thanks for the help how do I medal I am new to openstudy

OpenStudy (phi):

notice that tan x = sin(x)/cos(x) they may want you to answer with tan x

OpenStudy (anonymous):

sorry forgot to say that thanks anyway now how do I medal

OpenStudy (phi):

you click "best response" , like this

OpenStudy (anonymous):

thanks for the help

OpenStudy (phi):

yw